OpenStudy (anonymous):

y = 5x – 4 and 2y = 10x – 8 a) No solution b) One solution c) Infinitely many solutions

Directrix (directrix):

Which technique do you use to solve systems of equations? Some prefer substitution to Addition/Subtraction. Others don't.

OpenStudy (anonymous):

Whatever way is easier for you to explain. Doesnt matter

Directrix (directrix):

The site rules do not allow me to crank out the entire solution without your help. But, working together, we can knock this out. So, let's do it.

Directrix (directrix):

I look at the two equations and decide which technique will be easier to use. Looking at y = 5x – 4 and 2y = 10x – 8, I'll substitute (5x - 4) for y in the second equation 2y = 10x - 8. So, 2y = 10x - 8 becomes 2* ( 5x – 4 ) = 10 x - 8.

Directrix (directrix):

Your turn: Simplify this: 2* ( 5x – 4 ) = 10 x - 8. If you see how to continue to solve for x, do that. If not, stop and post the results after you simplify. I'd like to see your work.

Directrix (directrix):

@hanziebananzie